# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
$13^{\prime}$ $\overline{B}{}^{0}\to T^{+}_{c}K^{-}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{TP}-C^{\prime}_{PT,1})$
$13^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c}\pi^{-}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $B^{-}\to T^{\prime 0}_{c}K^{-}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}+3C^{\prime}_{TP}-2C^{\prime}_{PT,1})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}-C^{\prime}_{PT,1})$
$14^{\prime}$ $\overline{B}{}_{s}^{\prime 0}\to T^{\prime 0}_{c}\pi^{0}$ $-\frac{1}{2\sqrt{6}}(3C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\eta$ $\frac{\sqrt{3}c_{\phi^{\prime}}}{2\sqrt{2}}(-C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})+\frac{\sqrt{3}s_{\phi^{\prime}}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\eta^{\prime}$ $\frac{\sqrt{3}s_{\phi^{\prime}}}{2\sqrt{2}}(-C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})-\frac{\sqrt{3}c_{\phi^{\prime}}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c\bar{s}}K^{-}$ $-\frac{1}{2}(2C^{\prime}_{TP}-C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T})$
@@
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2\sqrt{2}}(C^{\prime}_{TP}+3E^{\prime}_{\bar{B}P,1}+4E^{\prime}_{\bar{B}P,2})-\frac{s_{\phi^{\prime}}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+4C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2\sqrt{2}}(C^{\prime}_{TP}+3E^{\prime}_{\bar{B}P,1}+4E^{\prime}_{\bar{B}P,2})+\frac{c_{\phi^{\prime}}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+4C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
@@
Note that the amplitudes of the following ten $\overline{B}\to TP$ modes in $\Delta S=0$ transitions in scenario II, namely $B^{-}\to T^{0}_{c\bar{s}}K^{-}$, $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\pi^{0}$, $\overline{B}{}^{0}\to T^{0}_{c\bar{s}}\overline{K}{}^{0}$, $\overline{B}{}^{0}\to T^{0}_{cs}K^{0}$, $B^{-}\to T^{-}_{cs}K^{0}$, $\overline{B}{}^{0}\to T^{-}_{cs}K^{+}$, $\overline{B}{}^{0}_{s}\to T^{+}_{c\bar{s}\bar{s}}K^{-}$, $B^{-}\to T^{-}_{c}\pi^{0}$, $\overline{B}{}^{0}\to T^{-}_{c}\pi^{+}$ and $\overline{B}{}_{s}^{0}\to T^{-}_{c}K^{+}$ decays, were also given in ref. Qin:2022nof . Five instead of eight flavor exotic states were considered. Their results agree with those shown in Table 13, but with $C_{T}$, $C_{PT}$, $C_{TP}$ and $E_{\bar{B}T}$ replaced with $T_{1}V_{cb}V^{*}_{ud}$, $T_{2}V_{cb}V^{*}_{ud}$, $T_{3}V_{cb}V^{*}_{ud}$ and $EV_{cb}V^{*}_{ud}$, respectively. Note that the rest in Table 13 and all in Tables 12 and 14 are new.
@@
$\Gamma(\overline{B}{}^{0}\to T^{0}_{c\bar{s}}\overline{K}{}^{0})$ $=$ $\left|\frac{V_{ud}}{V_{us}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to T^{0}_{cs}K^{0}),$ (63)
@@
We can compare the above relations to those in scenario I. We find that relations in Eqs. (63), (64) and (65) agree with those in Eqs. (59), (60) and (61), but those in Eq. (62) do not always agree with those in Eq. (58). In scenario I, as shown in Table 12, $B^{-}\to T^{0}_{c\bar{s}}K^{-}$ decay is governed by $C_{T}+C_{TP}$ and $B^{-}\to T^{0}_{cs}\pi^{-}$ and $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ decays are governed by $C^{\prime}_{T}+C^{\prime}_{TP}$. However, things are slightly different in scenario II. Although the first two modes are still governed by $C_{T}+C_{TP}$ or $C^{\prime}_{T}+C^{\prime}_{TP}$, the last one is governed by a different combination, $-C^{\prime}_{T}+C^{\prime}_{TP}$, see Tables 13 and 14. The discrepancy can be traced to the different isospin quantum numbers of $T^{0}_{cs}$ in these two scenarios. In scenario I, it is an isosinglet; there is only one way to match the isospin quantum numbers of the inintial state, $\overline{B}_{u,d}$, the operator $(\bar{c}b)(\bar{s}u)$ and the final states $T^{0}_{cs}$ $\pi$, while in scenario II, it is an isovector, and, hence, there is more than one way to match. Consequently, the $B^{-}\to T^{0}_{cs}\pi^{-}$ and $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ amplitudes can be rather different in the latter case.
@@
The above relations can be understood as in scenario I, but now the final states consist of $T$ in $\overline{\bf 15}$, with $P=K,\pi$, they form $\overline{\bf 15}\otimes{\bf 8}$. Furthermore, there are two $\overline{\bf 15}$, one ${\bf 6}$ and one ${\bf 3}$ in $\overline{\bf 15}\otimes{\bf 8}$, see Eq. (135), that can match the SU(3) quantum numbers of $\overline{B}$ and the operator. Although we have seven topological amplitudes, there are only four independent combinations, namely $C^{(\prime)}_{T}+C^{(\prime)}_{TP}$, $-C^{(\prime)}_{T}+C^{(\prime)}_{TP}$, $C^{(\prime)}_{T}+E^{(\prime)}_{\bar{B}P,1}$ and $C^{(\prime)}_{TP}-C^{(\prime)}_{PT,1}$, that contribute to modes with flavor exotic states. Note that the number also matches the number of topological amplitudes that can contribute to these modes. These modes are related.
@@ V.2.2 $\overline{B}\to TS$ decay amplitudes
$2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(C_{ST,1}+C_{ST,2}-C_{TS,1}-2C_{TS,2})$
@@
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}f_{0}$ $-\frac{c_{\phi}}{\sqrt{2}}(2C_{T}-C_{ST,1}-C_{ST,2}-C_{TS,1}-2C_{TS,2})$
$-s_{\phi}(C_{S}-C_{ST,1}-C_{TS,1})$
@@
6 $B^{-}\to T^{0}_{c}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(C_{S}-C_{T}+C_{ST,2}-C_{TS,1})$
6 $\overline{B}{}^{0}\to T^{0}_{c}a_{0}^{0}$ $-\frac{1}{2}(C_{T}-C_{ST,1}-C_{ST,2}+C_{TS,1}+2E_{\bar{B}S,1}+E_{\bar{B}T})$
6 $\overline{B}{}^{0}\to T^{0}_{c}f_{0}$ $-\frac{c_{\phi}}{2}(C_{T}-C_{ST,1}-C_{ST,2}-C_{TS,1}-2E_{\bar{B}S,1}-E_{\bar{B}T})$
$-\frac{s_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
7 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(2T_{S}+C_{S}+C_{ST,1}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}-E_{\bar{B}T})$
8 $B^{-}\to T^{\prime\prime 0}_{c}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(2T_{S}+C_{S}+C_{T}+C_{ST,2}+2C_{ST,3}+C_{TS,1})$
@@
$+\frac{s_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}+2C_{ST,3}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
8 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\sigma$ $-\frac{s_{\phi}}{2}(C_{T}-C_{ST,1}+C_{ST,2}+2C_{ST,3}-C_{TS,1}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+3E_{\bar{B}T})$
$+\frac{c_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}+2C_{ST,3}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
8 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\kappa^{0}$ $-\frac{1}{\sqrt{2}}(C_{T}+C_{S}-C_{ST,1}+2C_{ST,3}-2C_{TS,2})$
9 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c\bar{s}}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(C_{S}-C_{ST,2}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}-E_{\bar{B}T})$
9 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(2T_{S}+C_{ST,1}+C_{ST,2}+C_{TS,1}+2C_{TS,2})$
@@
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}a_{0}^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{S}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})$
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}f_{0}$ $-\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
$-s_{\phi}(C^{\prime}_{T}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$
@@
$-c_{\phi}(C^{\prime}_{T}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$

### v3

@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
$13^{\prime}$ $\overline{B}{}^{0}\to T^{+}_{c}K^{-}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{TP}-C^{\prime}_{PT,1})$
$13^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c}\pi^{-}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $B^{-}\to T^{\prime 0}_{c}K^{-}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}+3C^{\prime}_{TP}-2C^{\prime}_{PT,1})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}-C^{\prime}_{PT,1})$
$14^{\prime}$ $\overline{B}{}_{s}^{\prime 0}\to T^{\prime 0}_{c}\pi^{0}$ $\frac{1}{2\sqrt{6}}(3C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\eta$ $\frac{\sqrt{3}c_{\theta}}{2\sqrt{2}}(-C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})+\frac{\sqrt{3}s_{\theta}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\eta^{\prime}$ $\frac{\sqrt{3}s_{\theta}}{2\sqrt{2}}(-C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})-\frac{\sqrt{3}c_{\theta}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c\bar{s}}K^{-}$ $\frac{1}{2}(2C^{\prime}_{TP}-C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T})$
@@
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\theta}}{2\sqrt{2}}(C^{\prime}_{TP}+3E^{\prime}_{\bar{B}P,1}+4E^{\prime}_{\bar{B}P,2})-\frac{s_{\theta}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+4C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2\sqrt{2}}(C^{\prime}_{TP}+3E^{\prime}_{\bar{B}P,1}+4E^{\prime}_{\bar{B}P,2})+\frac{c_{\theta}}{2}(C^{\prime}_{T}+C^{\prime}_{PT,1}+4C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
@@
Note that the amplitudes of the following ten $\overline{B}\to TP$ modes in $\Delta S=0$ transitions in scenario II, namely $B^{-}\to T^{0}_{c\bar{s}}K^{-}$, $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\pi^{0}$, $\overline{B}{}^{0}\to T^{0}_{c\bar{s}}\overline{K}{}^{0}$, $\overline{B}{}^{0}\to T^{0}_{cs}K^{0}$, $B^{-}\to T^{-}_{cs}K^{0}$, $\overline{B}{}^{0}\to T^{-}_{cs}K^{+}$, $\overline{B}{}^{0}_{s}\to T^{+}_{c\bar{s}\bar{s}}K^{-}$, $B^{-}\to T^{-}_{c}\pi^{0}$, $\overline{B}{}^{0}\to T^{-}_{c}\pi^{+}$ and $\overline{B}{}_{s}^{0}\to T^{-}_{c}K^{+}$ decays, were also given in ref. Qin:2022nof . Five instead of eight flavor exotic states were considered. Their results agree with those shown in Table 13, but with $C_{T}$, $C_{PT,1}$, $C_{TP}$ and $E_{\bar{B}T}$ replaced with $T_{1}V_{cb}V^{*}_{ud}$, $T_{2}V_{cb}V^{*}_{ud}$, $T_{3}V_{cb}V^{*}_{ud}$ and $EV_{cb}V^{*}_{ud}$, respectively. Note that ref. Qin:2022nof also used a different notation for the open-charmed tetraquark states, see Table 2 for a comparison. For example, in Table 2 of ref. Qin:2022nof , one finds
$A(\overline{B}{}_{s}^{0}\to T_{cd\bar{u}\bar{s}}\pi^{0})=\frac{1}{\sqrt{2}}(T_{3}-T_{2})V_{cb}V^{*}_{ud},$ (91)
which corresponds to
$A(\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\pi^{0})=\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1}),$ (92)
in Table 13 but with $T^{0}_{c\bar{s}}$ replaced with $T_{cd\bar{u}\bar{s}}$, $C_{TP}$ replaced with $T_{3}V_{cb}V^{*}_{ud}$ and $C_{PT,1}$ replaced with $T_{2}V_{cb}V^{*}_{ud}$. Except for the above-mentioned modes, the rest in Table 13 and all in Tables 12 and 14 are new.
@@
[…]
@@
We can compare the above relations to those in scenario I. We find that relations in Eqs. (94), (95) and (96) agree with those in Eqs. (88), (89) and (90), but those in Eq. (93) do not always agree with those in Eq. (87). In scenario I, as shown in Table 12, $B^{-}\to T^{0}_{c\bar{s}}K^{-}$ decay is governed by $C_{T}+C_{TP}$ and $B^{-}\to T^{0}_{cs}\pi^{-}$ and $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ decays are governed by $C^{\prime}_{T}+C^{\prime}_{TP}$. However, things are slightly different in scenario II. Although the first two modes are still governed by $C_{T}+C_{TP}$ or $C^{\prime}_{T}+C^{\prime}_{TP}$, the last one is governed by a different combination, $C^{\prime}_{T}+C^{\prime}_{TP}$, see Tables 13 and 14. The discrepancy can be traced to the different isospin quantum numbers of $T^{0}_{cs}$ in these two scenarios. In scenario I, it is an isosinglet; there is only one way to match the isospin quantum numbers of the inintial state, $\overline{B}_{u,d}$, the operator $(\bar{c}b)(\bar{s}u)$ and the final states $T^{0}_{cs}$ $\pi$, while in scenario II, it is an isovector, and, hence, there is more than one way to match. Consequently, the $B^{-}\to T^{0}_{cs}\pi^{-}$ and $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ amplitudes can be rather different in the latter case.
@@
The above relations can be understood as in scenario I, but now the final states consist of $T$ in $\overline{\bf 15}$, with $P=K,\pi$, they form $\overline{\bf 15}\otimes{\bf 8}$. Furthermore, there are two $\overline{\bf 15}$, one $\bf 6$} and one $\bf 3$} in $\overline{\bf 15}\otimes{\bf 8}$, see Eq. (177), that can match the SU(3) quantum numbers of $\overline{B}$ and the operator. Although we have seven topological amplitudes, there are only four independent combinations, namely $C^{(\prime)}_{T}+C^{(\prime)}_{TP}$, $C^{(\prime)}_{T}+C^{(\prime)}_{TP}$, $C^{(\prime)}_{T}+E^{(\prime)}_{\bar{B}P,1}$ and $C^{(\prime)}_{TP}-C^{(\prime)}_{PT,1}$, that contribute to modes with flavor exotic states. Note that the number also matches the number of topological amplitudes that can contribute to these modes. These modes are related.
@@ V.2.2 $\overline{B}\to TS$ decay amplitudes
$2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(C_{ST,1}+C_{ST,2}-C_{TS,1}-2C_{TS,2})$
@@
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}f_{0}$ $\frac{c_{\phi}}{\sqrt{2}}(2C_{T}-C_{ST,1}-C_{ST,2}-C_{TS,1}-2C_{TS,2})$
$s_{\phi}(C_{S}-C_{ST,1}-C_{TS,1})$
@@
6 $B^{-}\to T^{0}_{c}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(C_{S}-C_{T}+C_{ST,2}-C_{TS,1})$
6 $\overline{B}{}^{0}\to T^{0}_{c}a_{0}^{0}$ $\frac{1}{2}(C_{T}-C_{ST,1}-C_{ST,2}+C_{TS,1}+2E_{\bar{B}S,1}+E_{\bar{B}T})$
6 $\overline{B}{}^{0}\to T^{0}_{c}f_{0}$ $\frac{c_{\phi}}{2}(C_{T}-C_{ST,1}-C_{ST,2}-C_{TS,1}-2E_{\bar{B}S,1}-E_{\bar{B}T})$
$\frac{s_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
7 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(2T_{S}+C_{S}+C_{ST,1}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}-E_{\bar{B}T})$
8 $B^{-}\to T^{\prime\prime 0}_{c}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(2T_{S}+C_{S}+C_{T}+C_{ST,2}+2C_{ST,3}+C_{TS,1})$
@@
$\frac{s_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}+2C_{ST,3}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
8 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\sigma$ $\frac{s_{\phi}}{2}(C_{T}-C_{ST,1}+C_{ST,2}+2C_{ST,3}-C_{TS,1}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+3E_{\bar{B}T})$
$\frac{c_{\phi}}{\sqrt{2}}(C_{S}+C_{T}-C_{ST,1}+2C_{ST,3}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}+E_{\bar{B}T})$
8 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}+C_{S}-C_{ST,1}+2C_{ST,3}-2C_{TS,2})$
9 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c\bar{s}}\kappa^{-}$ $\frac{1}{\sqrt{2}}(C_{S}-C_{ST,2}-C_{TS,1}-2C_{TS,2}+2E_{\bar{B}S,1}+2E_{\bar{B}S,2}-E_{\bar{B}T})$
9 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(2T_{S}+C_{ST,1}+C_{ST,2}+C_{TS,1}+2C_{TS,2})$
@@
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}a_{0}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{S}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})$
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}f_{0}$ $\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
$s_{\phi}(C^{\prime}_{T}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$
@@
$c_{\phi}(C^{\prime}_{T}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$